# Magnetic structure of $A\leq 10$ nuclei using the Norfolk nuclear models with quantum Monte Carlo methods

### v1

@@ II.3 The electromagnetic current operator
At N2LO, there is the two-body OPE current involving the excitation of an intermediate $\Delta$-isobar shown in panel (e) of Fig. 1. In the tables, we report the cumulative contribution at N2LO, that is the sum panels (d) and (e), and denote it with ‘N2LO’. When discussing the two-body magnetic densities, we analyze the N2LO OPE current involving the intermediate $\Delta$-isobar separately, and denote its contribution with ‘N2LO($\Delta$)’. This OPE term is purely IV and transverse, thus unconstrained by current conservation with the NV2 interaction.
@@ II.4 Magnetic form factors and moments
where $\lambda=\pm 1$, and $D_{M_{L},\lambda}^{L}$ are the rotation matrices [65]. The unit vector $\hat{\mathbf{e}}_{q\lambda}$ indicates the linear combination
@@ III.1.1 VMC magnetic moments of $A=3-7$ nuclei
Given that the $A=7$ nuclei are composed of two symmetric subgroups in the Young Scheme [72], having a dominant $[43]$ spatial symmetry pattern, there is an $\alpha+{}^{3}{\rm H}(^{3}{\rm He})$ cluster structure in ${}^{7}{\rm Li}~{}(^{7}{\rm Be})$. The nature of that clustering leads to the order-by-order breakdown of their magnetic moments being approximately the same as for ${}^{3}\rm H~{}(^{3}{\rm He})$. This is because the contribution to the magnetic moment within the $\alpha$ structure ($J_{\alpha}=0$) is small and any two-body contribution between the clusters is weak due the relatively weak OPE $N\!N$ strength in relative $P$ waves [72] suppressing meson-exchange contributions, and Pauli blocking suppressing inter-cluster contact contributions. This is made evident by the similarity of the two-body magnetic densities in Fig. 4, defined as
@@
Similarly, 6Li, with its dominant $[42]$ spatial symmetry, displays a strong $\alpha+d$ cluster structure. At leading order, the contribution to the magnetic moment in this system is approximately the value of the deuteron’s magnetic moment ($\mu_{d}$ = 0.857438 $\mu_{N}$ [73]), exhibiting minimal model dependence. Further, as 6Li is a purely IS ($T=0$) nucleus, the matrix elements for IV components of the current operator vanish. Consequently, some contributions from the current are suppressed beyond what one would expect from power counting considerations. For instance, the NLO contribution to the current is purely IV and thus does not contribute to the magnetic moment of 6Li. The N3LO(MIN) and N3LO(NM) contributions in this case are still enhanced, making few percent contributions to the overall magnetic moment. There is a ${\sim}15\%$ model dependence on the N3LO(MIN) contribution that is driven by changing the cutoff, whereas the N3LO(NM) contribution is much more stable across models.
@@ III.1.2 VMC magnetic moments of $A=8-10$ nuclei
Due to the cutoff dependence found in the one-pion range terms at NLO and N2LO, and the general model dependence observed in the N3LO terms, we focus on two representative model classes, Ia and IIb, for the heavier nuclei with $A=8-10$ to streamline computations.
@@
The addition of an unpaired $p$-shell neutron (proton) to ${}^{7}{\rm Li}~{}(^{7}{\rm Be})$ forms the predominantly [431] spatial symmetry ground state of ${}^{8}{\rm Li}~{}(^{8}{\rm B})$. At LO, the uncoupled valence nucleon quenches the magnitude of the magnetic moment. Interestingly, the NLO and N2LO contributions are not greatly altered by this additional nucleon. In these nuclei, there is little additional binding from the OPE $N\!N$ interaction with the added nucleon [72]. Consequently, the dominant one-pion range current contribution at NLO stays roughly the same in $A=8$ as in $A=7$ systems. The similarity is also true for the two-body density, as seen in Fig. 4. This carries over also into the long-range N2LO and N3LO contributions, while the short-range contributions at N3LO remain enhanced relative to their power counting, and are found to be model dependent. For 8B, two-body currents quench the magnetic moment by $\sim 30\%$. In 8Li, instead, they enhance it by $\sim 20\%$.
@@
For $A=9$ systems, the differences between the $T=3/2$ ground state of 9Li and the $T=1/2$ ground state of 9Be lead to interesting physics in the two-body currents’ effect. As discussed in the previous QMC evaluation within the hybrid approach [8], the [432] spatial symmetry dominates the 9Li ground state, giving it a $\alpha+t+2n$ cluster structure. On the other hand, 9Be is dominated by a [441] spatial symmetry and clusters as $2\alpha+n$. A single valence neutron outside of the $2\alpha$ cluster experiences almost no binding from the OPE $N\!N$ interaction [72]. Consequently, in this system, the two-body current contribution at NLO is suppressed relative to its power counting. This is reflected in the two-body densities of Fig. 6.
@@
The cluster structure also explains the suppression of the contact contributions in 9Be relative to 9Li. Pauli blocking suppresses contributions between the valence nucleon in 9Be and the $\alpha$ clusters. Instead, the two valence nucleons in the [432] cluster structure of 9Li form a spatially symmetric and IV pair. Thus, the long-range currents contributions are enhanced by this pair over the values for nuclei [43] and [431] cluster structures. The spatial symmetry also allows the additional pair of nucleons a non-zero overlap at shorter distances in their relative wave function. This explains the increased N3LO(NM) and N3LO(MIN) contribution in 9Li relative to 8Li.
@@ III.1.3 GFMC magnetic moments of $A=3-9$ nuclei
We note that two-body currents still play a large role in the total GFMC magnetic moments for all but the IS nuclei. As expected, the overall two-body correction still provides a small change to the LO result. Finally, in some cases, the two-body current contribution can change sign. These cases all correspond to currents whose two-body densities show the presence of a node due to an interplay of operators with opposite signs.
@@ III.1.4 Scaled two-body magnetic densities
The study of scaled two-body magnetic densities is relevant to identify universal behaviours in both short- and long-range two-nucleon dynamics. To examine the shape of the two-body magnetic densities, we scale them by the maximum value of $|r_{ij}^{2}\rho^{2b}_{M}|$, and analyze the scaled densities, defined as $r_{ij}^{2}\rho^{2b}_{M}/|r_{ij}^{2}\rho^{2b}_{M}|$. In Figs. 7 and 8, we plot the scaled IS and IV two-body magnetic densities, respectively, for all the nuclei considered in the present work.
@@ III.2 Magnetic form factors and high-momentum structure
The magnetic form factor has been computed using the same interactions and currents used for the magnetic moments. In Fig. 11, we present the results for all the nuclei showing the contributions of the currents order-by-order for the NV2+3-IIb interaction compared to data where available. We also include predicted magnetic form factors of radioactive isotopes, because while there are presently are no data for comparison, novel techniques have recently made it possible to study electron scattering from radioactive nuclei [84].
@@
For this analysis, we select a value of $\Lambda_{b}=700$ MeV and, for the characteristic scale of momentum $Q$, we follow Ref. [87] by setting it to the typical momentum transfer $(A-1)/A\,q$, to the nucleus of mass number $A$. In Fig. 13, we plot the error bands at LO (orange), NLO (green), N2LO (blue) and N3LO(red). Note that we do not renormalize the results to a specific value of the magnetic moment.
@@ III.2.6 Boron-10
The experimental data for the ${}^{10}\text{B}$ magnetic f.f. are provided for $0.5\lesssim q\lesssim 2.0$ fm-1 region [91, 100, 92, 98, 93]. We did not include the data of Refs. [91, 100] in the plots because of the strong disagreement with the more recent experimental results reported in Ref. [63]. The VMC calculation has a very nice agreement with the data in the region of the main peak (see Fig. 11l). Contributions coming from two-body currents are very small because they are purely IS. As shown in Fig. 12i, the main peak is dominated by the $M_{1}$ multiple, while the secondary peak in the $q\sim 2.5-3$ fm-1 region is generated by the $M_{5}$ multipole that also contributes to the filling in of the $M_{1}$’s diffraction minimum. Note that this secondary peak was hypothesized in Ref. [63], but it was not found in shell-model calculations performed in that reference. The $M_{3}$ peak that dominates all the other nuclei in the $q\sim 1.5$ fm-1 region is barely visible, confirming the findings of Ref. [63].
@@ IV Conclusions
In this work we presented the QMC calculations of magnetic moments and magnetic form factors of nuclei for $A\leq 10$ based on chiral effective field theory. For the calculation we used the Norfolk two- and three- body interaction and consistent one- and two-body current operators.
@@
To validate away from zero momentum transfer, we performed the calculation of magnetic form factors. In $A=3$ systems, our results represent a benchmark with Ref. [7] and we find good agreement with the previous evaluation and the data. We additionally performed the first microscopic calculations of magnetic form factors in $6\,\leq A\,\leq 10$ nuclei using a $\chi$EFT model, which we presented in a companion [49]. We observed good qualitative agreement with the data in most cases, reproducing the strength and position of peaks in the measured form factor. We also observed for the first time that, in the mirror nuclei, there is an inversion of the strength of the $M_{1}$ and $M_{3}$ multipoles. Here, we supplemented the discussion of Ref. [49] by demonstrating that this inversion is due to an interference between the LO orbital and spin-magnetization contributions to the $M_{1}$ multipole. As no experiments or previous calculations exist for these mirror systems, this represents a prediction of this effect.

### v2

@@ II.3 The electromagnetic current operator
At N2LO, there is the two-body OPE current involving the excitation of an intermediate $\Delta$-isobar shown in panel (e) of Fig. 1. In the tables, we report the cumulative contribution at N2LO, that is the sum of panels (d) and (e), and denote it with ‘N2LO’. When discussing the two-body magnetic densities, we analyze the N2LO OPE current involving the intermediate $\Delta$-isobar separately, and denote its contribution with ‘N2LO($\Delta$)’. This OPE term is purely IV and transverse, thus unconstrained by current conservation with the NV2 interaction.
@@ II.4 Magnetic form factors and moments
where $\lambda=\pm 1$, and $D_{M_{L},\lambda}^{L}$ are the Wigner rotation matrices [65]. The unit vector $\hat{\mathbf{e}}_{q\lambda}$ indicates the linear combination
@@ III.1.1 VMC magnetic moments of $A=3-7$ nuclei
Given that the $A=7$ nuclei are composed of two symmetric subgroups in the Young Scheme [72], having a dominant $\{43\}$ spatial symmetry pattern, there is an $\alpha+{}^{3}{\rm H}(^{3}{\rm He})$ cluster structure in ${}^{7}{\rm Li}~{}(^{7}{\rm Be})$. The nature of that clustering leads to the order-by-order breakdown of their magnetic moments being approximately the same as for ${}^{3}\rm H~{}(^{3}{\rm He})$. This is because the contribution to the magnetic moment within the $\alpha$ structure ($J_{\alpha}=0$) is small and any two-body contribution between the clusters is weak due the relatively weak OPE $N\!N$ strength in relative $P$ waves [72] suppressing meson-exchange contributions, and Pauli blocking suppressing inter-cluster contact contributions. This is made evident by the similarity of the two-body magnetic densities in Fig. 4, defined as
@@
Similarly, 6Li, with its dominant $\{42\}$ spatial symmetry, displays a strong $\alpha+d$ cluster structure. At leading order, the contribution to the magnetic moment in this system is approximately the value of the deuteron’s magnetic moment ($\mu_{d}$ = 0.857438 $\mu_{N}$ [73]), exhibiting minimal model dependence. Further, as 6Li is a purely IS ($T=0$) nucleus, the matrix elements for IV components of the current operator vanish. Consequently, some contributions from the current are suppressed beyond what one would expect from power counting considerations. For instance, the NLO contribution to the current is purely IV and thus does not contribute to the magnetic moment of 6Li. The N3LO(MIN) and N3LO(NM) contributions in this case are still enhanced, making few percent contributions to the overall magnetic moment. There is a ${\sim}15\%$ model dependence on the N3LO(MIN) contribution that is driven by changing the cutoff, whereas the N3LO(NM) contribution is much more stable across models.
@@ III.1.2 VMC magnetic moments of $A=8-10$ nuclei
Due to the strong cutoff dependence found in the one-pion range terms at NLO and N2LO, the lack of energy range dependence at those orders, and the general model dependence observed in the N3LO terms, we focus on two representative model classes, Ia and IIb, for the heavier nuclei with $A=8-10$ to streamline the computations.
@@
The addition of an unpaired $p$-shell neutron (proton) to ${}^{7}{\rm Li}~{}(^{7}{\rm Be})$ forms the predominantly $\{431\}$ spatial symmetry ground state of ${}^{8}{\rm Li}~{}(^{8}{\rm B})$. At LO, the uncoupled valence nucleon quenches the magnitude of the magnetic moment. Interestingly, the NLO and N2LO contributions are not greatly altered by this additional nucleon. In these nuclei, there is little additional binding from the OPE $N\!N$ interaction with the added nucleon [72]. Consequently, the dominant one-pion range current contribution at NLO stays roughly the same in $A=8$ as in $A=7$ systems. The similarity is also true for the two-body density, as seen in Fig. 4. This carries over also into the long-range N2LO and N3LO contributions, while the short-range contributions at N3LO remain enhanced relative to their power counting, and are found to be model dependent. For 8B, two-body currents quench the magnetic moment by $\sim 30\%$. In 8Li, instead, they enhance it by $\sim 20\%$.
@@
For $A=9$ systems, the differences between the $T=3/2$ ground state of 9Li and the $T=1/2$ ground state of 9Be lead to interesting physics in the two-body currents’ effect. As discussed in the previous QMC evaluation within the hybrid approach [8], the $\{432\}$ spatial symmetry dominates the 9Li ground state, giving it a $\alpha+t+2n$ cluster structure. On the other hand, 9Be is dominated by $\{441\}$ spatial symmetry and clusters as $2\alpha+n$. A single valence neutron outside of the $2\alpha$ cluster experiences almost no binding from the OPE $N\!N$ interaction [72]. Consequently, in this system, the two-body current contribution at NLO is suppressed relative to its power counting. This is reflected in the two-body densities of Fig. 6.
@@
The cluster structure also explains the suppression of the contact contributions in 9Be relative to 9Li. Pauli blocking suppresses contributions between the valence nucleon in 9Be and the $\alpha$ clusters. Instead, the two valence nucleons in the $\{432\}$ cluster structure of 9Li form a spatially symmetric and IV pair. Thus, the long-range currents contributions are enhanced by this pair over the values for nuclei with $\{43\}$ and $\{431\}$ cluster structures. The spatial symmetry also allows the additional pair of nucleons a non-zero overlap at shorter distances in their relative wave function. This explains the increased N3LO(NM) and N3LO(MIN) contribution in 9Li relative to 8Li.
@@ III.1.3 GFMC magnetic moments of $A=3-9$ nuclei
We note that two-body currents still play a large role in the total GFMC magnetic moments for all but the IS nuclei. As expected, the overall two-body correction still provides a small change to the LO result. In fact, recent many-body calculations of magnetic moments in heavy mass nuclei retaining only the first subleading correction at NLO have highlighted that two-body contributions remain important in heavy systems [80]. Finally, in some cases, the two-body current contribution can change sign. These cases all correspond to currents whose two-body densities show the presence of a node due to an interplay of operators with opposite signs.
@@ III.1.4 Scaled two-body magnetic densities
The study of scaled two-body magnetic densities is relevant to identify universal behaviours in both short- and long-range two-nucleon dynamics. Such universal behaviors can be understood in terms of generalized contacts [81] or from RG evolution [82]. In a similar spirit, to examine the shape of the two-body magnetic densities, we scale them by the maximum value of $|r_{ij}^{2}\rho^{2b}_{M}|$, and analyze the scaled densities, defined as $r_{ij}^{2}\rho^{2b}_{M}/|r_{ij}^{2}\rho^{2b}_{M}|$. In Figs. 7 and 8, we plot the scaled IS and IV two-body magnetic densities, respectively, for all the nuclei considered in the present work.
@@ III.2 Magnetic form factors and high-momentum structure
The magnetic form factor has been computed using the same interactions and currents used for the magnetic moments. In Fig. 11, we present the results for all the nuclei showing the contributions of the currents order-by-order for the NV2+3-IIb interaction compared to data where available. We also include predicted magnetic form factors of radioactive isotopes, because while there are presently no data for comparison, novel techniques have recently made it possible to study electron scattering from radioactive nuclei [86].
@@
For this analysis, we select a value of $\Lambda_{b}=700$ MeV and, for the characteristic scale of momentum $Q$, we follow Ref. [89] by setting it to the typical momentum transfer $(A-1)/A\,q$, for the nucleus of mass number $A$. In Fig. 13, we plot the error bands at LO (orange), NLO (green), N2LO (blue) and N3LO(red). Note that we do not renormalize the results to a specific value of the magnetic moment.
@@ III.2.6 Boron-10
The experimental data for the ${}^{10}\text{B}$ magnetic f.f. are provided for $0.5\lesssim q\lesssim 2.0$ fm-1 region [93, 102, 94, 100, 95]. We did not include the data of Refs. [93, 102] in the plots because of the strong disagreement with the more recent experimental results reported in Ref. [63]. The VMC calculation has a very nice agreement with the data in the region of the main peak (see Fig. 11l). Contributions coming from two-body currents are very small because they are purely IS. As shown in Fig. 12i, the main peak is dominated by the $M_{1}$ multiple, while the secondary peak in the $q\sim 2.5-3$ fm-1 region is generated by the $M_{5}$ multipole that also contributes to the filling in of the $M_{1}$’s diffraction minimum. Note that this secondary peak was hypothesized in Ref. [63], but it was not found in shell-model calculations performed in that reference. The $M_{3}$ peak that dominates all the other nuclei in the $q\sim 1.5$ fm-1 region is barely visible, as in Ref. [63].
@@ IV Conclusions
In this work we presented QMC calculations of magnetic moments and magnetic form factors of nuclei for $A\leq 10$ based on chiral effective field theory. For the calculation we used the Norfolk two- and three- body interaction and consistent one- and two-body current operators.
@@
To validate away from zero momentum transfer, we performed the calculation of magnetic form factors. In $A=3$ systems, our results represent a benchmark with Ref. [7] and we find good agreement with the previous evaluation and the data. We additionally performed the first microscopic calculations of magnetic form factors in $6\,\leq A\,\leq 10$ nuclei using a $\chi$EFT model, which we presented in a companion publication [49]. We observed good qualitative agreement with the data in most cases, reproducing the strength and position of peaks in the measured form factor. We also observed for the first time that, in the mirror nuclei, there is an inversion of the strength of the $M_{1}$ and $M_{3}$ multipoles. Here, we supplemented the discussion of Ref. [49] by demonstrating that this inversion is due to an interference between the LO orbital and spin-magnetization contributions to the $M_{1}$ multipole. As no experiments or previous calculations exist for these mirror systems, this represents a prediction of this effect.